When an object of mass 200g is submerged in methanol, its apparent weight is 1.052N. When submerged in benzene, its apparent weight is 0.951N. if the density of methanol is 0.8g/cm3, what is the density of benzene?
1
Expert's answer
2019-03-11T13:41:12-0400

According to the Archimedes law, an object immersed in a fluid is buoyed up by a force equal to the weight of the fluid displaced by the object. Thus, if the mass of the object is m and its volume is V, its apparent weight is mg ρVg = g(mρV), where ρ is the fluid density and g = 9.807 m/s2 is the acceleration of gravity. Denoting by ρb the density of benzene, and by ρm the density of methanol, we write the conditions of the problem as


g(mρmV) = Fm = 1.052 N , g(mρbV) = Fb = 0.951 N .


From these equations, we have, respectively,


ρmV = mFm/g , ρbV = mFb/g .


Dividing the second equation by the first one, we obtain


"\dfrac{\rho_{\rm b}}{\rho_{\rm m}} = \dfrac{m - F_{\rm b}/g}{m - F_{\rm m}/g}" or "\rho_{\rm b} = \dfrac{m - F_{\rm b}/g}{m - F_{\rm m}/g} \rho_{\rm m}" .


Substituting the numerical values, and taking into account that m = 0.2 kg in the International System of Units (SI), we have


"\rho_{\rm b} = \dfrac{ 0.2 - 0.951/9.807}{ 0.2 - 1.052/9.807} \rho_{\rm m} = 1.111 \rho_{\rm m} = 1.111 \times 0.8~ \text{g/cm}^{3 } = 0.889~\text{g/cm}^{3 }" .


Answer: 0.889 g/cm3.